OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 a. −2 b. −3 c. x can be any number as there are infinitely many solutions to the system. d. There is no x value as there is no solution to the system

OpenStudy (anonymous):

So we'll solve this problem using substitution. We are going to isolate x in the first equation okay. Once we do that 3x-6y=3 [Add 6y to both sides] 3x=6y+3 [Divide both sides by 3] x=2y+1.5 Now we put this into the 2nd equation and solve for y 7(2y+1.5)-5y=-11 Now you can solve the problem :) G2g now so yeah :(

OpenStudy (anonymous):

Thanks. <3